### v1

@@ Appendix I Lower bound
$||\mathcal{P}_{Yes}-\mathcal{P}_{Uniform}||_{1}$ $\leq||\mathcal{P}_{Yes}-\frac{\mathcal{P}_{Yes}}{\frac{19}{20}}||_{1}+||\frac{\mathcal{P}_{Yes}}{\frac{19}{20}}-\mathcal{P}_{Uniform}||_{1}$
$\leq\frac{1}{19}||\mathcal{P}_{Yes}||_{1}+\frac{\frac{1}{20}}{\frac{19}{20}}$
$=\frac{2}{19}$ since $\text{since }||\mathcal{P}_{Yes}||_{1}=1.$
@@
$||\mathcal{P}_{Yes}-\mathcal{P}_{Uniform}||_{TV}=\frac{1}{2}||\mathcal{P}_{Yes}-\mathcal{P}_{Uniform}||_{1}=\frac{1}{19}.$
Lemma I.8.
Let $\mathcal{D}_{Uniform}$ be the uniform distribution on the set of functions from $\mathbb{Z}_{p}^{Ct}$ to $\{-1,+1\}$. Then, the probability that the degree over $\mathbb{Z}_{p}$ of a randomly chosen function $f$ is $(2-\tau)$ close to $t$ in $\ell_{2}$ is $\leq\frac{1}{10}$.
Proof I.9 (Proof of Lemma I.8).
Let us fix a function $g$ from $\mathbb{Z}_{p}^{Ct}$ to $\{-1,+1\}$ such that $\deg_{p}(g)=t$. Let $f$ be randomly chosen function, where $f$ is also from $\mathbb{Z}_{p}^{Ct}$ to $\{-1,+1\}$. We need to show that the probability that the $\ell_{2}$ distance between $f$ and $g$ is $\leq$ $(2-\tau)$ is $\leq\frac{1}{10}$.
Let $\mathbb{I}_{x}$ be the indicator variable that takes the value $1$ when $f(x)\neq g(x)$, and it is 00 otherwise. Then $X_{g}^{\prime}=\frac{\sum_{x\in\mathbb{Z}_{p}^{Ct}}\mathbb{I}_{x}}{p^{Ct}}$ denotes the Hamming distance between them. Also, by Lemma A.15, if $X_{g}$ denotes thesquare of the $\ell_{2}$ distance between $f$ and $g$, then $X_{g}=4X_{g}^{\prime}$. Now,
$\mathbb{E}[X_{g}^{\prime}]=\frac{\sum_{x\in\mathbb{Z}_{p}^{Ct}}\mathbb{E}[\mathbb{I}_{x}]}{p^{Ct}}=\frac{1}{p^{Ct}}\frac{1}{2}p^{Ct}=\frac{1}{2}$
@@
Therefore, by Chernoff (Lemma A.21),
@@
$=\Pr[|\sum_{x\in\mathbb{Z}_{p}^{Ct}}\mathbb{I}_{x}-\mathbb{E}[\sum_{x\in\mathbb{Z}_{p}^{Ct}}\mathbb{I}_{x}]|\geq\frac{\tau}{4}\times p^{Ct}]$
$\leq 2\exp(-\frac{2}{3}\bigl{(}\frac{\tau}{4}\bigr{)}^{2}p^{Ct}),$
@@
$\Pr[X_{g}\leq 2-\tau]\leq 2\exp(-\frac{1}{24}\tau^{2}p^{Ct}).$
Now, the number of functions with degree over $\mathbb{Z}_{p}$ equal to $t$ is $\leq p^{\binom{Ct}{t}}$. The probability that $f$ is $(2-\tau)$-close to a degree $t$ function in $\ell_{2}$ is given by,
$\Pr[\cup_{g:\deg(g)=t}\{X_{g}\leq 2-\tau\}]$
$\leq p^{\binom{Ct}{t}}\times 2\exp(-\frac{2}{3}\tau^{2}p^{Ct})$
$\leq 2\exp\biggl{(}(aC)^{t}\ln p-\frac{2}{3}\tau^{2}p^{Ct}\biggr{)}$ by Stirling’s approximation
@@
taking $C\geq\frac{1}{t\ln p}\ln\biggl{(}\frac{3((aC)^{t}\ln p+\ln 20)}{2\tau^{2}}\biggr{)}$, where $a$ is a constant.
Corollary I.10.
$||\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}$.
Proof I.11.
Let $A$ be the set of all functions $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ which are $(2-\tau)$ close to $\mathbb{Z}_{p}$-degree $t$ in $\ell_{2}$. Then $\mathcal{P}_{No}(A)=0$, and by Lemma I.8, $\mathcal{P}_{Uniform(A)}\leq\frac{1}{10}$, where $\mathcal{P}_{No}$ and $\mathcal{P}_{Uniform}$ denote the probability over the distributions $\mathcal{D}_{No}$ and $\mathcal{D}_{Uniform}$ respectively. So, $|\mathcal{L}_{No}(A)-\mathcal{L}_{Uniform}(A)|\leq\frac{1}{10}$.
@@
$||\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}.$
Proof I.12 (Proof of Theorem I.2).
From Lemma I.6, we have $||\mathcal{L}_{Yes}-\mathcal{L}_{Uniform}||\leq\frac{1}{19}$. Also, from Corollary I.10, we have $||\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}$. Therefore, by triangle inequality,
$||\mathcal{L}_{Yes}-\mathcal{L}_{No}||_{TV}$
$\leq||\mathcal{L}_{Yes}-\mathcal{L}_{Uniform}||_{TV}+||\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}$
@@
$||\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}=||\mathcal{L}_{Yes}-\mathcal{L}_{No}||_{TV}\leq\frac{1}{3}.$
Remark I.13.
The proof of Theorem I.2 (hence Theorem 1.9) holds for $\mathcal{G}$, since $\deg_{p}$ of a Boolean valued function over $\mathbb{Z}_{p}^{n}$ can be defined for a Boolean valued function over $\mathcal{G}$ in a similar manner, and the proof goes along the lines of the proof of Theorem I.2.

### v2

@@ 8 Query lower bound on adaptively testing sparsity
[…]
@@
[…]
Lemma 8.6.
Let $\mathcal{D}_{Uniform}$ be the uniform distribution on the set of functions from $\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}$ to $\{-1,+1\}$. Then, the probability that the degree over $\mathcal{G}$ of a randomly chosen function $f$ is $(2-\tau)$ close to $t$ in $\ell_{2}$ is $\leq\frac{1}{10}$.
Proof of Lemma 8.6.
Let us fix a function $g$ from $\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}$ to $\{-1,+1\}$ such that $\deg_{\mathcal{G}}(g)=t$. Let $f$ be randomly chosen function, where $f$ is also from $\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}$ to $\{-1,+1\}$. We need to show that the probability that the $\ell_{2}$ distance between $f$ and $g$ is $\leq$ $(2-\tau)$ is $\leq\frac{1}{10}$.
Let $\mathbb{I}_{x}$ be the indicator variable that takes the value 11 when $f(x)\neq g(x)$, and it is 0 otherwise. Then $X_{g}^{\prime}=\frac{\sum_{x\in\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}}\mathbb{I}_{x}}{p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}}$ denotes the Hamming distance between them. Also, by Lemma 2.43, if $X_{g}$ denotes thesquare of the $\ell_{2}$ distance between $f$ and $g$, then $X_{g}=4X_{g}^{\prime}$. Now,
$\mathbb{E}[X_{g}^{\prime}]=\frac{\sum_{x\in\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}}\mathbb{E}[\mathbb{I}_{x}]}{p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}}=\frac{1}{p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}}\frac{1}{2}p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}=\frac{1}{2}$
@@
Therefore, by Chernoff (Lemma 2.30),
@@
$=\Pr[|\sum_{x\in\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}}\mathbb{I}_{x}-\mathbb{E}[\sum_{x\in\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}}\mathbb{I}_{x}]|\geq\frac{\tau}{4}\times p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}]$
$\leq 2\exp(-\frac{2}{3}\bigl{(}\frac{\tau}{4}\bigr{)}^{2}p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}),$
@@
$\Pr[X_{g}\leq 2-\tau]\leq 2\exp(-\frac{1}{24}\tau^{2}p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}).$
Now, the number of functions with degree over $\mathcal{G}$ equal to $t$ is $\leq n^{\prime\binom{(C+1)t}{t}}$, where $n^{\prime}=\operatorname{LCM}\{p_{i_{1}}^{m_{i_{1}}},\ldots,p_{i_{Ct}}^{m_{i_{Ct}}}\}$.
The probability that $f$ is $(2-\tau)$-close to a degree $t$ function in $\ell_{2}$ is given by,
$\Pr[\cup_{g:\deg_{\mathcal{G}}(g)=t}\{X_{g}\leq 2-\tau\}]$
$\leq n^{\prime\binom{(C+1)t}{t}}\times 2\exp(-\frac{2}{3}\tau^{2}p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}})$
$=2\exp\biggl{(}\binom{(C+1)t}{t}\ln n^{\prime}-\ln t!-\frac{2}{3}\tau^{2}p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}\biggr{)}$
$\leq 2\exp\biggl{(}\binom{(C+1)t}{t}\ln n^{\prime}-\ln t!-\frac{2}{3}\tau^{2}(\min\{p_{i_{1}^{m_{i_{1}}}},\ldots,p_{i_{Ct}^{n_{i_{Ct}}}}\})^{Ct}\biggr{)}$
@@
taking $C\geq\frac{1}{t\ln\min\{p_{i_{1}^{m_{i_{1}}}},\ldots,p_{i_{Ct}^{m_{i_{Ct}}}}\}}\ln\biggl{(}\frac{3(\binom{(C+1)t}{t}\ln n^{\prime}-\ln t!+\ln 20)}{2\tau^{2}}\biggr{)}-1$. ∎
Corollary 8.7.
$\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}$.
Proof.
Let AA be the set of all functions $f:\mathbb{Z}_{p_{i_{1}}^{n_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{n_{i_{Ct}}}}\to\{-1,+1\}$ which are $(2-\tau)$ close to $\deg_{\mathcal{G}}=t$ in $\ell_{2}$. Then $\mathcal{P}_{No}(A)=0$, and by Lemma 8.6, $\mathcal{P}_{Uniform(A)}\leq\frac{1}{10}$, where $\mathcal{P}_{No}$ and $\mathcal{P}_{Uniform}$ denote the probability over the distributions $\mathcal{D}_{No}$ and $\mathcal{D}_{Uniform}$ respectively. So, $|\mathcal{L}_{No}(A)-\mathcal{L}_{Uniform}(A)|\leq\frac{1}{10}$.
@@
$\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}.$
Proof of Theorem 8.2.
From Lemma 8.5, we have $\mathcal{L}_{Yes}-\mathcal{L}_{Uniform}||\leq\frac{1}{19}$. Also, from Corollary 8.7, we have $\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}$. Therefore, by triangle inequality,
[…]
@@
[…]
9 Conclusion
Gopalan et al. [GOS+11] was the first to study the problem of testing Fourier sparsity of Boolean function over $\mathbb{Z}^{n}_{2}$. Along the way, they were able to drive fundamental properties of Boolean functions over $\mathbb{Z}_{2}^{n}$, like Granularity of the Fourier spectrum, that have found many other applications [ACL+21b, CMM+21]. In this work, we have extended their results for finite Abelian groups.
Finally, we ask whether it is possible to show a better ($\Delta$-dependent) lower bound over $\mathcal{G}$ on the query-complexity of any adaptive sparsity testing algorithm.